### v3

@@ 4 Frequentist Regret Analysis of POFUL
In this section, we present the frequentist regret analysis of POFUL algorithms. We first introduce useful concentration events that hold with high-probability.
@@
Fix $\delta\in(0,1)$ and $\delta^{\prime}=\delta/2T$. We define $\beta^{PVT}_{t}(\delta^{\prime})\coloneqq\iota_{t}(\delta)\beta^{RLS}_{t}(\delta^{\prime})$ and
$\ \widehat{\mathcal{A}}_{t}$$\coloneqq\left\{\forall s\leq t:\left\|\widehat{\theta}_{t}-\theta^{\star}\right\|_{V_{t}}\leq\beta^{RLS}_{t}(\delta^{\prime})\right\},\quad\widetilde{\mathcal{A}}_{t}\coloneqq\left\{\forall s\leq t:\left\|\widetilde{\theta}_{t}-\widehat{\theta}_{t}\right\|_{V_{t}}\leq\beta^{PVT}_{t}(\delta^{\prime})\right\}.$
We also define $\mathcal{A}_{t}\coloneqq\widehat{\mathcal{A}}_{t}\cap\widetilde{\mathcal{A}}_{t}$
@@
In the following, we decompose the regret of POFUL into three terms, $\mathcal{R}^{PE}(T)$, $\mathcal{R}^{PVT}(T)$ and $\mathcal{R}^{RLS}(T)$, and sketch how they are bounded separately. Let $x_{t}^{\star}=\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\theta^{\star}}\rangle$ be the benchmark action and $\widetilde{x}_{t}$ be the action chosen by POFUL, the regret decomposes as
$\mathcal{R}(T)$ $=\sum_{t=1}^{T}\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle+\sum_{t=1}^{T}\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle+\sum_{t=1}^{T}\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle$
$\eqqcolon\mathcal{R}^{PE}(T)+\mathcal{R}^{PVT}(T)+\mathcal{R}^{RLS}(T).$
$\mathcal{R}^{RLS}(T)$ is the regret due to the estimation error of the RLS estimator. Note that $\widetilde{x}_{t}$’s are the action sequence used to construct the RLS estimate. By Proposition 2 their cumulative $V_{t}^{-1}$-norm is bounded by $2d\log(1+t/\lambda)$. Hence the upper bound of $\mathcal{R}^{RLS}(T)$ is essentially determined by the $V_{t}$-distance between $\widehat{\theta}_{t}$ and $\theta^{\star}$, which are characterized by the confidence ellipsoids $\mathcal{E}^{RLS}_{t}$. $\mathcal{R}^{PVT}(T)$ corresponds to the regret due to the exploration of POFUL by choosing the pivot parameter $\widetilde{\theta}_{t}\in\mathcal{E}^{PVT}_{t}$. Similar to $\mathcal{R}^{RLS}(T)$, the upperbound is related to the $V_{t}$-distance between $\widetilde{\theta}_{t}$ and $\widehat{\theta}_{t}$, which is controlled by construction and depends on the inflation parameter $\iota_{t}$. As a result, it can be shown that $\mathcal{R}^{RLS}(T)=\widetilde{\mathcal{O}}(d\sqrt{T})$ and $\mathcal{R}^{PVT}(T)=\widetilde{\mathcal{O}}(d(\sum_{t=1}^{T}\iota_{t}^{2})^{1/2})$ with probability $1-\delta$. Notably, $\mathcal{R}^{RLS}(T)=\widetilde{\mathcal{O}}(d\sqrt{T})$, matching the known minimax lowerbound. For $\mathcal{R}^{PVT}(T)$ to match this lower bound as well, one way is to set $\iota_{t}=\widetilde{\mathcal{O}}(1)$ for all $t\in[T]$. However, later we see this could be problematic for bounding $\mathcal{R}^{PE}(T)$.
$\mathcal{R}^{PE}(T)$ corresponds to the pessimism term. In the Bayesian analysis of LinTS, the expectation of this term is 00, with respect to $\theta^{\star}$, as long as $\widetilde{\theta}_{t}$ and $\theta^{\star}$ are sampled from the same distribution which occurs when LinTS has access to the true prior distribution for $\theta^{\star}$. In the frequentist analysis, however, we need to control the pessimism term incurred by any random sample $\widetilde{\theta}_{t}$. For OFUL, this term is bounded properly with high probability by the construction of OFUL actions. For LinTS, the only known analysis due to Agrawal and Goyal (2013); Abeille et al. (2017) gives a bound of order $\widetilde{\mathcal{O}}(d\sqrt{dT})$ using an optimism-based approach, which is worse than the Bayesian regret by a factor of $\sqrt{d}$. The key component of their proof is introducing inflation to the posterior variance by setting $\iota_{t}=\widetilde{\mathcal{O}}(\sqrt{d})$. In Abeille et al. (2017), the inflation is necessary for them to lower bound the probability that LinTS samples optimistic $\widetilde{\theta}_{t}$ that leads to a nonpositive pessimism i.e. $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle\leq\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle$ with a constant probability. For Greedy, there isn’t a general theoretical guarantee to bound the pessimism term without additional structures. The way of bounding it distinguishes our method from the methods based on optimism (Agrawal and Goyal, 2013; Abeille et al., 2017), and enables us to improve the frequentist regret of LinTS and Greedy that cannot be analyzed by the optimism-based methods.
In fact, non-optimistic samples might be able to control the pessimism term. Our method does not force choosing optimistic samples and hence there is no need to inflate the posterior. As a result, it does not introduce a $\sqrt{d}$-gap in the regret. Instead, we introduce a measurement of the “quality” of POFUL actions with respect to the regret, and develop a computable upper bound for this quality using geometric approaches. The main merits are an instance-dependent regret bound for POFUL that matches the minimax optimal regret in some settings, and variants of LinTS and Greedy with provable frequentist regret bound that is minimax optimal up to only a constant factor.
4.1 An Instance-Dependent Regret Bound for POFUL
In the following, we condition on the event $\mathcal{A}_{T}$ which holds with probability $1-\delta$. The following proposition bounds the instantaneous regret of POFUL and we defer the proof to Appendix 10.2.
@@ 4.1 An Instance-Dependent Regret Bound for POFUL
Suppose ${\theta}^{\star}\in\mathcal{E}_{t}^{RLS}(\delta^{\prime})$ and $\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})$, it holds that
$\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime}).$
On the right-hand side, since the oracle optimal action sequence $\{x^{\star}_{t}\}_{t\in[T]}$ is unknown to the algorithm and is different from the action sequence $\{\widetilde{x}_{t}\}_{t\in[T]}$ played by POFUL, one cannot apply Proposition 2 to bound the summation $\sum_{t=1}^{T}\left\|\widetilde{x}_{t}\right\|_{V_{t}^{-1}}^{2}$ to get an upperbound of the regret. To fix the problem due to this discrepancy, the key point is connecting $\{\widetilde{x}_{t}\}_{t\in[T]}$ and $\{x^{\star}_{t}\}_{t\in[T]}$ in terms of the $V_{t}^{-1}$-norm. This motivates the following definition.
@@
Note that $\langle{x},{\widehat{\theta}_{t}-\theta^{\star}}\rangle\leq\|x\|_{V_{t}^{-1}}\beta_{t}^{RLS}$ holds with high probability, we have that $\|x\|_{V_{t}^{-1}}$ essentially determines the length of the confidence interval of the reward $\langle{x},{\theta^{\star}}\rangle$. Hence, $\alpha_{t}$ serves as the ratio of uncertainty degrees of the reward obtained by the optimal action $x^{\star}_{t}$ and the chosen action $\widetilde{x}_{t}$.
The intuition behind the definition for $\mu_{t}$ is constructing a regret upper bound similar to that of OFUL. Specifically, Proposition 4 indicates $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq\mu_{t}\|\widetilde{x}_{t}\|_{V_{t}^{-1}}\beta^{RLS}_{t}$, and we can check that the instantaneous regret of OFUL satisfies $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq 2\|\widetilde{x}_{t}\|_{V_{t}^{-1}}\beta^{RLS}_{t}$. In this sense, $\mu_{t}$ is a proxy of the instantaneous regret incurred by POFUL at time $t$. Moreover, OFUL can be regarded as a POFUL algorithm whose $\mu_{t}$ is fixed at 2, and we could extend the definition of $\alpha_{t}$ to OFUL by solving $\mu_{t}=\alpha_{t}(1+\iota_{t}-\tau_{t})+1+\iota_{t}+\tau_{t}$ and set $\alpha_{t}\equiv 1$ for OFUL (recall $\iota_{t}\equiv 0$ and $\tau_{t}\equiv 1$ for OFUL).
The following Theorem connects $\{\mu_{t}\}_{t\in[T]}$ and $\mathcal{R}(T)$ and we postpone the proof to Appendix 10.3. It provides an oracle but general frequentist regret upper bound for all POFUL algorithms.
@@
Fix $\delta\in(0,1)$ and let $\delta^{\prime}=\delta/2T$. Under Assumptions 1, 2 and 3, with probability $1-\delta$, POFUL achieves a frequentist regret of
$\mathcal{R}(T)\leq\sqrt{2d\left(\sum_{t=1}^{T}\mu_{t}^{2}\right)\log\left(1+\frac{T}{\lambda}\right)}\beta_{T}^{RLS}(\delta^{\prime}).$ (4)
@@
We call Theorem 1 an oracle regret bound as $\{\mu_{t}\}_{t\in[T]}$ for general POFUL depends on the unknown system parameter $\theta^{\star}$. In general, they cannot be calculated by the decision-maker. Nevertheless, note that $(1+\iota_{t}-\tau_{t})\geq 0$ for OFUL, LinTS, and Greedy. Suppose we can derive computable upper bounds $\{\widehat{\alpha}_{t}\}_{t\in[T]}$ for $\{\alpha_{t}\}_{t\in[T]}$ respectively, then we could calculate upper bounds for $\{\mu_{t}\}_{t\in[T]}$ as well, and Theorem 1 instantly turns into a data-driven regret bound for POFUL, and could be utilized later for the course correction. This is indeed our aim in the next section and we achieve it using geometric arguments.
@@
A similar notion of (the reciprocal of) our $\alpha_{t}$ can be found in the Discussion section in Abeille et al. (2017), where they indicate their proof of regret bound for LinTS can be reproduced if LinTS samples $\widetilde{\theta}_{t}$ satisfying $\|x^{\star}(\widetilde{\theta}_{t})\|_{V_{t}^{-1}}\geq\alpha\|x^{\star}(\theta^{\star}_{t})\|_{V_{t}^{-1}}$ with constant probability for some $\alpha>0$. They comment that the set of these parameters doesn’t coincide with the set of optimistic parameters and it remains an open problem whether one could relax the requirement of oversampling the posterior through this observation. Our work answers this question positively in the next section by investigating the property of (the reciprocal of their) $\alpha$ using geometric arguments, providing an explanation for LinTS’s empirical success.
@@ 5 A Data-Driven Approach
Typically, we can check $\widehat{\alpha}_{t}$ is no less than one because we might have $\widetilde{x}_{t}=x_{t}^{\star}$. As a result, it holds that $\widehat{\mu}_{t}=(1+\widehat{\alpha}_{t})(1+\iota_{t})+(1-\widehat{\alpha}_{t})\tau_{t}\geq(1+\widehat{\alpha}_{t})(1+\iota_{t})$, i.e., non-zero $\tau_{t}$ gives a larger regret proxy. Given this observation, we’ll focus on scenarios where $\tau_{t}\equiv 0$. This avoids the large regret proxies and computational inefficiencies seen in OFUL when determining the action $\widetilde{x}_{t}$. Such scenarios include LinTS and its variants like TS-Freq as well as Greedy - all of which are standard algorithms that still lack theoretical regret guarantee results.
In the following, we construct such upper bounds $\{\widehat{\alpha}_{t}\}_{t\in[T]}$ for both discrete and continuous action scenarios.
5.1 Discrete Action Space
We start with the discrete action space scenario to illustrate our method. It turns out that upper bounding $\alpha_{t}$ is done by simply comparing the confidence interval lengths of all potentially optimal actions.
$L_{t}(x)\coloneqq\langle{x},{\widehat{\theta}_{t}}\rangle-\iota_{t}\|x\|_{V_{t}^{-1}}\beta^{RLS}_{t}(\delta^{\prime})\leq\langle{x},{{\theta}^{\star}}\rangle\leq\langle{x},{\widehat{\theta}_{t}}\rangle+\iota_{t}\|x\|_{V_{t}^{-1}}\beta^{RLS}_{t}(\delta^{\prime})\eqqcolon U_{t}(x).$
the set of potentially optimal actions is given by $C_{t}(\beta_{t}^{RLS}(\delta^{\prime}))$. Similarly, the set of actions that might be chosen by POFUL is given by $\mathcal{C}_{t}(\beta_{t}^{PVT}(\delta^{\prime}))$, where we let $\beta^{PVT}_{t}\coloneqq\iota_{t}\beta^{RLS}_{t}$. The inflation parameter $\iota_{t}$ turns out to control the conservativeness when eliminating actions.
$\widehat{\alpha}_{t}\coloneqq\frac{\sup_{x\in\mathcal{C}_{t}(\beta_{t}^{RLS}(\delta^{\prime}))}\|x\|_{V_{t}^{-1}}}{\inf_{x\in\mathcal{C}_{t}(\beta_{t}^{PVT}(\delta^{\prime}))}\|x\|_{V_{t}^{-1}}}.$ (5)
5.2 Continuous action space.
When it comes to the generic settings where the number of actions is large or when the set of actions is a continuous set in $\mathbb{R}^{d}$, the upper bound above is inefficient or even infeasible to compute. In this section, we tackle this issue. Our strategy capitalizes on geometric insights related to the properties of the confidence ellipsoids, providing upper bounds that can be computed efficiently.
@@ 5.1 Discrete Action Space
For the sake of a better illustration, we consider $\mathcal{X}_{t}\equiv{\mathcal{S}}_{d-1}$ for this scenario, which is a standard example of continuous action space, and is the same as the setting considered in Abeille et al. (2017). We remark that for this specific setting, the problem is still hard. This is because we don’t have a closed-form solution for the set of potentially optimal actions.
@@
In this setting, the optimal action $x_{t}^{\star}(\theta)\coloneqq\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\theta}\rangle$ takes a closed form $x^{\star}_{t}(\theta)=\theta/\|\theta\|$. To upper bound $\alpha_{t}$, a direct idea is to consider respectively the smallest and largest value of $\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}$ for $\theta$ in the confidence ellipsoids of $\theta$, namely, $\mathcal{E}^{RLS}_{t}$ and $\mathcal{E}^{PVT}_{t}$. Specifically, we have
@@
$\alpha_{t}\leq\frac{\sup_{\theta\in\mathcal{E}_{t}^{RLS}}\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}}{\inf_{\theta\in\mathcal{E}_{t}^{PVT}}\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}}.$ (6)
@@
As is illustrated in Figure 1. The set of potentially optimal actions $\mathcal{C}_{t}$ is the projection of the confidence ellipsoid $\mathcal{E}_{t}$ onto ${\mathcal{S}}_{d-1}$. It’s hard to get a closed-form expression for $\mathcal{C}_{t}$, so we cannot directly calculate the range of $V_{t}^{-1}$-norm of actions in $\mathcal{C}_{t}$. Nevertheless, we can approximate the range by investigating the geometric properties of the ellipsoids.
The intuition here is that, when POFUL has implemented sufficient exploration so that $\mathcal{E}_{t}$ is small enough, $\mathcal{C}_{t}$ concentrates accordingly to a small cap on ${\mathcal{S}}_{d-1}$. All actions within $\mathcal{C}_{t}$ point to similar directions and thus have similar $V_{t}$-norm. (Note that for a unit vector $x$, we have that $\|x\|_{V_{t}}$ can be written as a weighted summation of the eigenvalues of $V_{t}$, where the weights are determined by the direction of $x$.) Hence we can bound the range of $V_{t}$-norm by investigating the concentration of $\mathcal{C}_{t}$, which is pure geometry, and we’ll use the range of $V_{t}$-norm to determine the range of $V_{t}^{-1}$-norm.
@@
Figure 1: Illustration of the set $\mathcal{C}_{t}$ of potentially optimal actions in $\mathbb{R}^{2}$. (a): $\mathcal{C}_{t}$ is the projection of $\mathcal{E}_{t}$ onto ${\mathcal{S}}_{d-1}$. (b): As the decision-maker collects more and more data, the confidence ellipsoid $\mathcal{E}_{t}$ shrinks accordingly (here different color corresponds to different levels of exploration). The potentially optimal actions point to similar directions, which essentially determine their $V_{t}$-norm. Hence it’s plausible one could get a proper range of their $V_{t}$-norm using geometry.
@@
The following main theorem of this paper derives an upper bound of $\alpha_{t}$ based on this idea. We sketch the proof in Section 5.3.

### v5

@@ 4 Frequentist Regret Analysis of POFUL
In this section, we present the frequentist regret analysis of POFUL algorithms. Proofs appear in Appendix 11. We first introduce high-probability concentration events.
@@
Fix $\delta\in(0,1)$ and $\delta^{\prime}=\delta/2T$. We define $\beta^{PVT}_{t,\delta^{\prime},\lambda_{\text{reg}}}\coloneqq\iota_{t}\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}$, $\widehat{\mathcal{A}}_{t}\coloneqq\{\forall s\leq t:\|\widehat{\theta}_{t}-\theta^{\star}\|_{V_{t}}\leq\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}\},\quad\widetilde{\mathcal{A}}_{t}\coloneqq\{\forall s\leq t:\|\widetilde{\theta}_{t}-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta^{PVT}_{t,\delta^{\prime},\lambda_{\text{reg}}}\}$, and $\mathcal{A}_{t}\coloneqq\widehat{\mathcal{A}}_{t}\cap\widetilde{\mathcal{A}}_{t}$.
@@
4.1 An Data-Driven Regret Bound for POFUL
In the following, we condition on the event $\mathcal{A}_{T}$ which holds with probability $1-\delta$. The following proposition bounds the instantaneous regret of POFUL.
@@ 4.1 An Data-Driven Regret Bound for POFUL
Suppose ${\theta}^{\star}\in\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}$ and $\widetilde{\theta}_{t}\in\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{PVT}$, it holds that
$\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}.$ (4)
Note that this upper bound is different from what’s used in the optimism-based methods (Abbasi-Yadkori et al., 2011; Agrawal and Goyal, 2013; Abeille et al., 2017), we reproduce their upper bound and discuss the relationship of our method and theirs in Appendix 12.
On the right-hand side of (4), since the oracle optimal action sequence $\{x^{\star}_{t}\}_{t\in[T]}$ is unknown to the algorithm and is different from the action sequence $\{\widetilde{x}_{t}\}_{t\in[T]}$ played by POFUL, one cannot apply Proposition 2 to bound the summation $\sum_{t=1}^{T}\|\widetilde{x}_{t}\|_{V_{t}^{-1}}^{2}$ and get an upperbound of the regret. To address this, the key point to connect $\{\widetilde{x}_{t}\}_{t\in[T]}$ and $\{x^{\star}_{t}\}_{t\in[T]}$ with $V_{t}^{-1}$-norm. This motivates the following definition.
@@
Note that $\langle{x},{\widehat{\theta}_{t}-\theta^{\star}}\rangle\leq\|x\|_{V_{t}^{-1}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}$ holds with high probability, we have that $\|x\|_{V_{t}^{-1}}$ essentially determines the length of the confidence interval of the reward $\langle{x},{\theta^{\star}}\rangle$. Hence, $\alpha_{t}$ serves as the ratio of uncertainty degrees of the reward obtained by the optimal action $x^{\star}_{t}$ and the chosen action $\widetilde{x}_{t}$.
The intuition behind the definition for $\mu_{t}$ is constructing a regret upper bound similar to that of OFUL. Specifically, Proposition 4 indicates $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq\mu_{t}\|\widetilde{x}_{t}\|_{V_{t}^{-1}}\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}$, and we can check that the instantaneous regret of OFUL satisfies $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq 2\|\widetilde{x}_{t}\|_{V_{t}^{-1}}\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}$. In this sense, $\mu_{t}$ is a proxy of the instantaneous regret incurred by POFUL at time $t$. Moreover, OFUL can be regarded as a POFUL algorithm whose $\mu_{t}$ is fixed at 2, and we could extend the definition of $\alpha_{t}$ to OFUL by solving $\mu_{t}=\alpha_{t}(1+\iota_{t}-\tau_{t})+1+\iota_{t}+\tau_{t}$ and set $\alpha_{t}=1$ for all $t\in[T]$ for OFUL (recall that in OFUL, $\iota_{t}=0$ and $\tau_{t}=1$ for all $t\in[T]$).
The following Theorem connects $\{\mu_{t}\}_{t\in[T]}$ and $\mathcal{R}(T)$. It provides an oracle but general frequentist regret upper bound for all POFUL algorithms.
@@
Fix $\delta\in(0,1)$ and let $\delta^{\prime}=\delta/2T$. Under Assumptions 1, 2 and 3, with probability $1-\delta$, POFUL achieves a regret of
$\mathcal{R}(T)\leq\sqrt{2d\left(\sum_{t=1}^{T}\mu_{t}^{2}\right)\log\left(1+\frac{T}{\lambda_{\text{reg}}}\right)}\beta_{T,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}\,.$ (5)
@@
We call Theorem 1 an oracle regret bound as $\{\mu_{t}\}_{t\in[T]}$ for general POFUL depends on the unknown system parameter $\theta^{\star}$. In general, they cannot be calculated by the decision-maker. Nevertheless, note that $\iota_{t}$ and $\tau_{t}$ are chosen by the decision-maker, when we have computable upper bounds $\{\widehat{\alpha}_{t}\}_{t\in[T]}$ for $\{\alpha_{t}\}_{t\in[T]}$, using $\mu_{t}^{2}\leq 2\alpha_{t}^{2}(1+\iota_{t}-\tau_{t})^{2}+2(1+\iota_{t}+\tau_{t})^{2}$, we could calculate upper bounds for $\{\mu_{t}\}_{t\in[T]}$ as well. Consequently, Theorem 1 instantly turns into a data-driven regret bound for POFUL and could be utilized later for course correction, which will be the aim of the next section. When we additionally know that $1+\iota_{t}-\tau_{t}$ is non-negative, we would use the equality $\mu_{t}=\alpha_{t}(1+\iota_{t}-\tau_{t})+1+\iota_{t}+\tau_{t}$ directly for the bound.
@@
In the Discussion section of Abeille et al. (2017), the authors introduce a concept similar to the reciprocal of our $\alpha_{t}$. They suggest that the necessity of proving LinTS samples are optimistic could be bypassed if for some $\alpha>0$ LinTS samples $\widetilde{\theta}_{t}$ such that $\|x^{\star}(\widetilde{\theta}_{t})\|_{V_{t}^{-1}}\geq\alpha\|x^{\star}(\theta^{\star}_{t})\|_{V_{t}^{-1}}$ with constant probability, where $x^{\star}(\widetilde{\theta}_{t})$ and $x^{\star}(\theta^{\star}_{t})$ represent the optimal actions corresponding to $\widetilde{\theta}_{t}$ and $\theta^{\star}_{t}$, respectively. They pose this as an open question regarding the possibility of relaxing the requirement of inflating the posterior. In the following section, we provide a positive answer to this question by studying the reciprocal of their $\alpha$ using geometric arguments. This investigation offers an explanation for the empirical success of LinTS without the need for posterior inflation.
@@ 5 A Data-Driven Approach
We focus on scenarios where $\tau_{t}=0$ for all $t\in[T]$. These include LinTS and variants such as TS-Freq, as well as Greedy - standard algorithms still lacking theoretical regret guarantees. Below, we construct upper bounds $\{\widehat{\alpha}_{t}\}_{t\in[T]}$ for the continuous-action scenario. Bounds for discrete-action scenarios appear in Appendix 13.
5.1 Continuous action space
@@
Our strategy capitalizes on geometric insights related to the properties of the confidence ellipsoids, providing upper bounds that can be computed efficiently. For the sake of a better illustration, we consider $\mathcal{X}_{t}={\mathcal{S}}_{d-1}$ for all $t\in[T]$ for this scenario, where ${\mathcal{S}}_{d-1}=\left\{x\in\mathbb{R}^{d}:\|x\|=1\right\}$ is the unit hypersphere in $\mathbb{R}^{d}$. This is a standard example of continuous action space, and is the same as the setting considered in Abeille et al. (2017). We remark that for this specific setting, the problem is still hard. This is because we don’t have a closed-form solution for the set of potentially optimal actions.
@@
In this setting, the optimal action $x_{t}^{\star}(\theta)\coloneqq\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\theta}\rangle$ takes the form $x^{\star}_{t}(\theta)=\theta/\|\theta\|$. To upper bound $\alpha_{t}$, we consider respectively the smallest and largest value of $\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}$ for $\theta$ in the confidence ellipsoids of $\theta$, namely, $\mathcal{E}^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}$ and $\mathcal{E}^{PVT}_{t,\delta^{\prime},\lambda_{\text{reg}}}$. Specifically, we have
@@
$\alpha_{t}\leq\frac{\sup_{\theta\in\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}}\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}}{\inf_{\theta\in\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{PVT}}\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}}.$ (6)
@@
As is illustrated in Figure 2, the set of potentially optimal actions $\mathcal{C}_{t}$ is the projection of the confidence ellipsoid $\mathcal{E}_{t}$ onto ${\mathcal{S}}_{d-1}$. It’s hard to get a closed-form expression for $\mathcal{C}_{t}$, so we cannot directly calculate the range of $V_{t}^{-1}$-norm of actions in $\mathcal{C}_{t}$. Nevertheless, when POFUL has implemented sufficient exploration so that $\mathcal{E}_{t}$ is small enough, $\mathcal{C}_{t}$ concentrates accordingly to a small cap on ${\mathcal{S}}_{d-1}$. Therefore, it is possible to estimate the range of the $V_{t}$-norm by employing geometric reasoning. Subsequently, this estimated range will be utilized to ascertain the range of the $V_{t}^{-1}$-norm.
@@
Figure 2: Illustration of potentially optimal actions set $\mathcal{C}_{t}$ in $\mathbb{R}^{2}$. (a): $\mathcal{C}_{t}$ is $\mathcal{E}_{t}$’s projection onto ${\mathcal{S}}_{d-1}$. (b): As more data is collected, $\mathcal{E}_{t}$ shrinks (colors show exploration levels). Potentially optimal actions point in similar directions, determining their $V_{t}$-norm. This suggests their $V_{t}$-norm range could be estimated geometrically.
@@
The main theorem (proved in Section 9) derives an upper bound for $\alpha_{t}$ based on this idea.